### v1

@@ Abstract
In recent years, Multimodal Large Language Models (MLLMs) have increasingly emphasized grounding and referring capabilities to achieve detailed understanding and flexible user interaction. However, in the realm of visual document understanding, these capabilities lag behind due to the scarcity of fine-grained datasets and comprehensive benchmarks. To fill this gap, we propose the DOcument Grounding and rEferring data engine (DOGE-Engine), which produces two types of high-quality fine-grained document data: multi-granular parsing data for enhancing fundamental text localization and recognition capabilities; and instruction-tuning data to activate MLLM’s grounding and referring capabilities during dialogue and reasoning. Additionally, using our engine, we construct DOGE-Bench, which encompasses 7 grounding and referring tasks across 3 document types (chart, poster, PDF document), providing comprehensive evaluations for fine-grained document understanding. Furthermore, leveraging the data generated by our engine, we develop a strong baseline model, DOGE. This pioneering MLLM is capable of accurately referring and grounding texts at multiple granularities within document images. Our code, data, and model will be open-sourced for community development.
@@
Figure 1: We propose DOGE, a multi-modal large language model that enables users to engage in versatile grounded document interactions. The figure illustrates the reasoning results of DOGE for grounding, grounding-and-referring, and referring tasks in DOGE-Bench.
@@ 1 Introduction
In recent years, Multimodal Large Language Models (MLLMs) [14, 33, 63, 64, 62, 32, 26, 21, 6, 23, 25] have achieved significant advancements in general visual understanding and reasoning by integrating pre-trained vision encoders with large language models. Leveraging the diverse fine-grained annotations of existing image datasets, some researchers further equip MLLMs with grounding and referring capabilities. These capabilities enhance the detailed visual understanding, increase the credibility of responses, and facilitate more efficient human-AI interaction.
In the field of visual document understanding, the dense textual content and complex layout significantly complicate fine-grained understanding.
To build a user-friendly and trustworthy document AI assistant, it is crucial to enable users to refer to specific regions of a document for more precise comprehension and provide accurate grounding of key details for more expressive and effective interaction.
However, due to the lack of high-quality, fine-grained document datasets, the full potential of MLLMs in document grounding and referring remains largely unexplored. Existing efforts leverage the multi-granularity parsing annotations to enhance document detailed perception [12], or introduce region-level instruction-tuning tasks to achieve basic referring capabilities [28]. However, these works have two significant shortcomings:
@@
    Suboptimal annotation quality. To collect parsing annotations, the Optical Character Recognition (OCR) tools are typically employed to extract text and bounding boxes. Subsequently, random document regions and overlapping text boxes are selected to create multi-granular annotations. However, OCR tools have issues with inaccurate bounding box predictions, and the random selection can lead to truncated characters and imprecise bounding boxes. Furthermore, the text annotated using this approach often suffers from incomplete semantics, limiting its potential for application in developing instruction-tuning data.
@@
    Lack of diversity in task formats. Current instruction-tuning datasets primarily cater to fundamental referencing tasks, such as region-level OCR and summarization. However, these datasets fall short in supporting grounding tasks and fail to seamlessly incorporate grounding and referencing capabilities into the dialogue and reasoning processes of MLLMs. The limited range of tasks hinders the model’s ability to perceive details flexibly and impacts the user interaction experience.
@@
In order to advance the document referring and grounding capability, we propose the DOcument Grounding and rEferring data engine (DOGE-Engine) for high-quality fine-grained document data construction. With DOGE-Engine, we create 1) 1.4M multi-granular document parsing data, which includes text box annotations at the word, phrase, line, paragraph and full-page level across poster, chart and PDF document. This dataset is utilized to enhance basic text localization and recognition capabilities and serves as the foundation for creating instruction-tuning data; 2) a diverse set of 700K instruction-tuning data. This includes both text-in location-out (grounding) and location-in text-out (referring) data, as well as data that combines location and text in both input and output. These instruction-tuning data are constructed based on our multi-granular document parsing data via the assistance of GPT-4o [14], possessing high linguistic quality and accurate ground-and-refer annotations.
@@
Furthermore, in visual document understanding, evaluation tasks related to grounding and referring are relatively scarce, making it challenging to assess the model’s corresponding capabilities. To this end, we propose DOGE-Bench, which contains 4K test samples and encompasses 7 grounding and referring tasks across 3 document types (chart, poster, PDF document). Finally, based on data generated by our engine, we develop a strong baseline model, DOGE, capable of understanding spatial referring and accurately grounding text within document images. We report the performance of our model on DOGE-Bench, providing a performance reference for future research.
@@
In summary, our contributions are threefold. (1) We introduce DOGE-Engine, a data construction pipeline that generates large-scale, high-quality multi-granular document parsing data and diverse ground-and-refer instruction tuning data. (2) We develop DOGE-Bench, the first comprehensive benchmark designed to evaluate MLLM’s grounding and referring capabilities in document understanding. (3) We present DOGE, a pioneering MLLM that is capable of understanding referred text and performing text grounding during conversation.
@@
Figure 2: Left: Hierarchical task taxonomy and sample distribution analysis. Right: The pipeline of DOGE-Engine, which outlines the construction process for multi-granularity parsing data and ground-and-refer instruction-tuning data.
@@ 2.1 MLLMs for Visual Document Understanding
Visual Document Understanding focuses on comprehending images with extensive text content. Recently, several Multimodal Large Language Models [61, 56, 9, 6] have been introduced to perform visual document understanding without relying on OCR tools. UReader [61] unifies a wide range of document understanding tasks with instruction-tuning format. A shape-adaptive cropping module is further designed to encode rich textual content in high-resolution image, where the raw image is cropped into multiple tiles, with each tile being individually encoded to represent features of the raw image. TextMonkey [34] employs shifted window to build connections between different image tiles, alleviating the issue of incoherence semantic caused by image cropping. InternLM-XComposer-4KHD [9] and InternVL 1.5 [7] further increase the tile number, significantly improving the performance on visual document understanding tasks. These works achieve promising performance but lack document grounding and referring capabilities, which hinders the grounded document understanding and flexible human-AI interaction.
@@ 2.2 MLLMs for Grounding and Referring
In pursuit of fine-grained image understanding and convenient interaction, recent studies integrate grounding and referring abilities into MLLMs [4, 44, 65, 68]. Kosmos2 [44], Shikra [4] and Ferret [65] utilize bounding boxes or visual prompts to pinpoint specific regions of an image and generate responses with key objects being grounded, facilitating flexible content referring and interaction. Additionally, LLaVA-Grounding [68] and GLaMM [45] further employ finer-grained multi-granularity masks for pixel-level grounding across various semantic levels. These works achieve promising grounding and referring results on real-world images but cannot be adapted to visual document understanding due to the image domain gap. Recently, there are several attempts to develop visual document grounding and referring. mPLUG-1.5 [12] and Kosmos-2.5[35] enhance localization and fine-grained perception with bounding box annotations from OCR tools and PDF parsing tools, respectively. Fox [29] utilizes various visual prompts to refer to document regions and enables the MLLMs to extract or translate the region-level content. However, mPLUG-1.5 and Kosmos-2.5 only support basic text localization and recognition tasks. Fox further supports the referring translation task.
These methods, though effective, fall short of integrating both grounding and referring into the broader reasoning and dialogue processes. As a result, they leave the full potential of MLLMs for visual document grounding and referring largely untapped, particularly in reasoning tasks where grounding plays a crucial role.
3 DOGE-Engine
Well-annotated and diverse grounded data are crucial for improving the grounding and referring capabilities of MLLMs [44, 52, 67, 27]. Currently, there is a shortage of comprehensive and accurately labeled document grounded data. Manually annotating raw document images is both time-consuming and labor-intensive, as it requires not only marking the bounding boxes but also accurately annotating all the text within those boxes. To tackle this challenge, we collect a substantial volume of documents and develop the DOGE-Engine to construct fine-grained document grounded datasets. Our document data sources primarily encompass three document data types: posters, charts, and PDF documents. As shown in the right of Fig. 4, we start by filtering the raw data to remove low-quality samples or those with missing or broken information. Then, we introduce the generation of two types of high-quality document data: The annotation process of multi-granular parsing data is detailed in Sec. 3.1, and the construction pipeline of instruction-tuning data is elaborated in Sec. 3.2. An overview of statistical information is provided in Sec. 3.3. Annotation results are presented in Appendix A.
@@ 3.1.1 Automatic Bounding box Annotation
Figure 3: (a) For the poster and chart data, we propose the Re-rendering Strategy to automatically obtain precise bounding boxes. (b) For the PDF document data, we propose the Merge Strategy combining the annotations from MinerU and PyMuPDF to obtain comprehensive and layout-aware full-page parsing annotations
Poster. We collect poster data from the Crello dataset [59], which consists of design templates from the design service website. Each poster contains the document meta-annotation, which includes rendered text blocks and their corresponding bounding boxes. However, we observe that some provided rendered text blocks are significantly degraded, and the bounding boxes are inaccurate. To address this issue, we re-render the text content and design a Re-rendering Strategy to obtain precise bounding boxes. Specifically, as illustrated in Fig. 3(a), the strategy consists of three steps: Firstly, we render the poster with the meta-annotation; Then, we modify the color or opacity attribute of one text block and perform a re-rendering process; Finally, since the rendering results of first two steps are identical except for the modified attribute, the image of the target text block can be obtained by applying pixel-wise subtraction. By repeating the above steps for every text block, we can obtain accurate bounding boxes of all text blocks, which are then normalized to create the final annotation. Note that we preserve the origin font format during re-rendering to keep the font diversity of poster data.
@@
PDF Document. CC-MAIN-2021-31-PDF-UNTRUNCA TED consists of a large collection of various text documents in PDF format. We select a high-quality subset of moderately sized files. Leveraging the PDF parsing tool PyMuPDF and a document content extraction framework MinerU[55], we directly extract parsing data of different granularities (word, phrase, line, paragraph).
@@ 3.1.2 Full-page Parsing Data Construction
Poster and Chart. The poster data features less text, making it straightforward to grasp the dependencies between text blocks. Therefore, we concatenate the paragraph-level text and their corresponding bounding boxes in a left-to-right, top-to-bottom scanning order to create the full-page parsing annotation in JSON format. Due to the highly structured nature of chart data, its full-page parsing annotation is organized as a JSON dictionary containing key-value pairs of chart, axis labels, legends, and their corresponding bounding boxes.
PDF Document. PDF document contains extensive text and complex layouts, necessitating parsing in a logical reading order to comprehend the information accurately. MinerU contains a layout detection model and an OCR model, allowing for extracting text blocks with a certain reading order. However, it often fails to capture all content in documents, missing tables, footnotes, and other elements. In contrast, PyMuPDF can thoroughly extract content from documents, but it does not provide an appropriate reading order. We propose a Merge Strategy that combines the strengths of MinerU and PyMuPDF to achieve comprehensive, layout-aware annotations. As shown in Fig. 3(b), 1) we first compare text blocks from ordered and unordered maps to eliminate duplicate text blocks. For the truncated blocks (red block 1 and red block 2 in the ordered map), we replace them with the corresponding complete block (blue block in the unordered map) to improve the semantic completeness within blocks; 2) In the ordered map, we construct an ordered area if two consecutive blocks are placed from the top left to the bottom right. We classify the preserved blocks in the unordered map into two categories: in-area blocks and out-of-area blocks; 3) We insert the in-area blocks into the ordered area and sequentially update the block order in each ordered area using column-major order; 4) For the out-of-area blocks, we insert them into the ordered map. The order of these blocks is then determined based on their positional relationship with their nearest ordered blocks, following column-major order.
@@ 3.2 Instruction-Tuning Data Construction
To seamlessly integrate grounding and referring capabilities in the dialogue and reasoning, it is crucial to construct high-quality instruction-tuning data with accurate grounded text in both query and response. Based on the precise bounding boxes annotations from Sec. 3.1, we leverage GPT-4o [14] to generate diverse-formatted instruction-tuning data tailored to various text granularities, encompassing tasks such as question answering, summarization and reasoning. During construction, we also constrain the length and type of the generated response, resulting in different response categories, e.g., short response, open-ended long response. Subsequently, we add a response format prompt to each query to improve the instruction-following quality of generated data. Prompts for GPT-4o are shown in Appendix B
Poster and Chart. As introduced in Sec. 3.1.2, the full-page parsing annotation for poster and chart data includes all text blocks along with their corresponding bounding boxes, which contain sufficient information to construct ground-and-refer instruction-tuning data. Therefore, we provide GPT-4o with the full-page parsing data, and task GPT-4o to generate queries and responses in a grounded manner, where the generated content must include texts that originate from the given parsing annotations. Additionally, we require GPT-4o to warp the texts originating from parsing annotations with ‘<ocr></ocr>’ and append the corresponding coordinates wrapped in ‘<bbox></bbox>’.
PDF Document. PDF document often contains extensive text content. If we input all the text of various granularities into GPT-4o, it would result in excessively long prompts. This not only incurs significant API calling costs but also increases the difficulty for GPT-4o in comprehending and completing the tasks, leading to low generation quality. To address these issues, we design a Post-annotating Strategy to generate high-quality data with minimal overhead. Specifically, 1) We first use the document image as input for GPT-4o instead of the text of document, which significantly reduces the token count; 2) Then, we task GPT-4o to generate queries and responses based on the content of document images. Additionally, any text within the queries and responses that originates from the document image must be wrapped with “<ocr></ocr>”; 3) We extract the texts wrapped with “<ocr></ocr>” and utilize the PyMuPDF to retrieve the corresponding bounding boxes; 4) Finally, we warp these bounding boxes with ‘<bbox></bbox>’ and insert them back into the generated content after their corresponding texts. For texts that cannot be located by PyMuPDF, we remove wrapped tokens and convert them to plain text. This approach allows the generation of diverse instruction-tuning data at a low cost while ensuring the quality of grounding and referring annotations.

### v2

@@ Abstract
With recent advances in Multimodal Large Language Models (MLLMs), grounding and referring capabilities have gained increasing attention for achieving detailed understanding and flexible user interaction. However, these capabilities still remain underdeveloped in visual document understanding due to the scarcity of fine-grained datasets and comprehensive benchmarks. To fill this gap, we propose the DOcument Grounding and Referring data engine (DOGR-Engine), which generates two types of high-quality fine-grained document data: (1) multi-granular parsing data to improve text localization and recognition, and (2) instruction-tuning data to activate MLLMs’ grounding and referring capabilities in dialogue and reasoning. Using the DOGR-Engine, we construct DOGR-Bench, a benchmark covering seven grounding and referring tasks across three document types (chart, poster, and PDF document), offering a comprehensive evaluation of fine-grained document understanding. Leveraging the generated data, we further develop DOGR, a strong baseline model that excels in text localization and recognition, while precisely grounds and refers to key textual information during conversation and reasoning, thereby advancing document understanding to a finer granularity and enable flexible interaction paradigms. Our code, data, and model are open-sourced at https://github.com/zyinan99/DOGR.
@@
Figure 1: We propose DOGR, a multi-modal large language model that enables users to engage in versatile grounded document interactions. The figure illustrates the reasoning results of DOGR for grounding, grounding-and-referring, and referring tasks in DOGR-Bench.
@@ 1 Introduction
Visual document understanding is challenging due to dense textual content and complex layouts, which hinder fine-grained comprehension. Effective interaction with such documents requires the ability to refer to specific regions for precise understanding and accurately ground key details.
However, due to the lack of a clear definition for document grounding and referring tasks, as well as the absence of high-quality fine-grained data for training and evaluation, the advancement of multimodal large language models in this field has been notably limited. Many existing models [1, 7, 16, 47] struggle with basic fine-grained document tasks such as text localization and region-level recognition, let alone achieve effective grounding and referring to crucial textual elements during conversation or reasoning.
Several prior works either incorporate multi-granularity parsing data during pre-training to enhance document perception [14] or introduce simple region-level instruction-tuning tasks to achieve basic referring capabilities [24]. While these methods aim to improve document understanding, they all suffer from two significant shortcomings:
@@
    Suboptimal annotation quality. To collect parsing annotations, the Optical Character Recognition (OCR) tools are typically employed to extract text and bounding boxes. However, OCR tools encounter issues with inaccurate recognition of text and corresponding bounding boxes and have difficulty extracting semantically coherent text from documents with complex layouts, limiting their potential for developing fine-grained document understanding and question answering data.
@@
    Lack of diversity in task formats. Current instruction-tuning datasets primarily cater to basic referring tasks, such as region-level OCR, summarization and translation. However, these datasets fall short in supporting grounding tasks and fail to seamlessly incorporate grounding and referencing capabilities into the dialogue and reasoning processes of MLLMs. The restricted range of tasks hinders the model’s ability to flexibly perceive details and affects the user interaction experience.
@@
To solve the above two issues, we introduce the DOcument Grounding and rEferring data engine (DOGR-Engine) for constructing high-quality fine-grained document data. DOGR-Engine generates: (1) 2.1M multi-granular document parsing data, which provides text box annotations at the word, phrase, line, paragraph and full-page level across posters, charts, and PDF documents. This dataset enhances text localization and recognition capabilities and serves as the foundation for creating instruction-tuning data. (2) A diverse set of 700K instruction-tuning samples, covering text-in location-out (grounding), location-in text-out (referring) data, and tasks combining location and text in both input and output. These data, generated from our multi-granular document parsing dataset with GPT-4o [16] assistance, ensure high linguistic quality and accurate grounding and referring annotations.
To benchmark the MLLMs’ document grounding and referring capabilities comprehensively, as well as providing a clear task definition,
we propose DOGR-Bench, which contains 3.6K test samples and encompasses seven grounding and referring tasks across three document types (chart, poster, PDF document). On this benchmark, the state-of-the-art MLLMs demonstrate poor performance for grounding and referring tasks, suggesting that current models may not yet be equipped to handle the intricate challenges in fine-grained document understanding. This performance gap underscores the DOGR-Bench’s value in identifying key limitations and guiding future advancements in MLLMs.
@@
Furthermore, based on data generated by our engine, we develop a strong baseline model, DOGR, capable of understanding spatial referring and accurately grounding text within document images. We report the performance of our model on DOGR-Bench, providing a performance reference for future research.
@@
In summary, our contributions are threefold. (1) We introduce DOGR-Engine, a data construction pipeline that generates large-scale, high-quality multi-granular document parsing data and diverse ground-and-refer instruction tuning data. (2) We develop DOGR-Bench, the first comprehensive benchmark designed to evaluate MLLMs’ grounding and referring capabilities in document understanding. (3) We present DOGR, a pioneering MLLM that is capable of understanding referred text and performing text grounding during conversation, leading to detailed document comprehension and more intuitive user interaction.
@@
Figure 2: Left: Hierarchical task taxonomy and sample distribution analysis. Right: The pipeline of DOGR-Engine, which outlines the construction process for multi-granularity parsing data and ground-and-refer instruction-tuning data.
@@ 2.1 MLLMs for Visual Document Understanding
Recently, several Multimodal Large Language Models [59, 54, 12, 6, 66, 22] have been introduced to perform visual document understanding without relying on OCR tools. UReader [59] unifies a wide range of document understanding tasks with instruction-tuning format. A shape-adaptive cropping module is further designed to encode rich textual content in high-resolution image. TextMonkey [29] employs shifted window to build connections between different image tiles, alleviating the issue of incoherence semantic caused by image cropping. InternLM-XComposer-4KHD [12] and InternVL 1.5 [8] further increase the tile number, significantly improving the performance on visual document understanding tasks. These works achieve promising performance but lack document grounding and referring capabilities, which hinders the grounded document understanding and flexible human-AI interaction.
@@ 2.2 MLLMs for Grounding and Referring
In pursuit of fine-grained image understanding and convenient interaction, recent studies integrate grounding and referring abilities into MLLMs [4, 39, 60, 63, 1, 58]. Kosmos2 [39], Shikra [4] and Ferret [60] utilize bounding boxes or visual prompts to pinpoint specific regions of an image and generate responses with key objects being grounded, facilitating flexible content referring and interaction. Additionally, LLaVA-Grounding [63] and GLaMM [40] further employ finer-grained multi-granularity masks for pixel-level grounding across various semantic levels. These works perform well on real-world images but can’t adapt to visual document understanding due to domain gaps. Recently, there are several attempts to develop visual document grounding and referring. mPLUG-1.5 [14] and Kosmos-2.5[30] enhance localization and fine-grained perception with bounding box annotations from OCR tools and PDF parsing tools, respectively. Fox [25] utilizes various visual prompts to refer to document regions and enables the MLLMs to extract or translate the region-level content. However, mPLUG-1.5 and Kosmos-2.5 only support basic text localization and recognition tasks. Fox further supports the referring translation task. These methods, though effective, fall short of integrating both grounding and referring into broader reasonings and dialogues. As a result, they leave the full potential of MLLMs for visual document grounding and referring largely untapped, particularly in reasoning tasks where grounding plays a crucial role.
3 DOGR-Engine
While well-annotated and diverse grounded data are crucial for improving the grounding and referring abilities of MLLMs [39, 49, 62, 23], comprehensive and accurately labeled document-grounded datasets still remain scarce. Given the time-consuming and labor-intensive nature of manually annotating raw document images, we collect a large corpus of documents, including posters, charts, and PDF files, and develop DOGR-Engine to efficiently construct fine-grained document-grounded datasets.
In this section, we first introduce the data processing pipeline and the generation of two types of high-quality document data. Specifically, the data processing pipeline and the annotation process for multi-granular parsing data are described in Sec. 3.1, while the construction pipeline for instruction-tuning data is elaborated in Sec. 3.2. Additionally, an overview of statistical information is provided in Sec. 3.3, and the annotation results are in Appendix A.
@@ 3.1 Multi-granular Parsing Data Construction
As shown in Fig. 2(Right), we start by filtering the raw data to remove low-quality samples or those with missing or broken information. We then describe the annotation process for multi-granular parsing data as follows.
@@ 3.1.1 Automatic Bounding box Annotation
Figure 3: (a) For the poster and chart data, we propose the Re-rendering Strategy to automatically obtain precise bounding boxes. (b) For the PDF document data, we propose the Merge Strategy combining the annotations from MinerU and PyMuPDF to obtain comprehensive and layout-aware full-page parsing annotations.
Poster. We collect poster data from the Crello dataset [57], which consists of poster templates from the design service website. Each poster contains the document meta-annotation, which includes rendered text blocks and their corresponding bounding boxes. However, we observe that some provided rendered text blocks are significantly degraded, and the bounding boxes are inaccurate. To address this issue, we re-render the text content and design a Re-rendering Strategy to obtain precise bounding boxes. Specifically, as illustrated in Fig. 3(a), the strategy consists of three steps: Firstly, we render the poster with the meta-annotation; Then, we modify the color or opacity attribute of one text block and perform a re-rendering process; Finally, since the rendering results of the first two steps are identical except for the modified attribute, the image of the target text block can be obtained by applying pixel-wise subtraction. By repeating the above steps for every text block, we can obtain accurate bounding boxes of all text blocks, which are then normalized to create the final annotation. Note that we preserve the original font format during re-rendering to keep the font diversity of poster data.
@@
PDF Document. We select a high-quality subset from CC-MAIN-2021-31-PDF-UNTRUNCATED [48], which consists of a large collection of various text documents in PDF format. Leveraging the PDF parsing tool PyMuPDF [11] and the document content extraction framework MinerU [52], we directly extract parsing data of different granularities (word, phrase, line, paragraph).
@@ 3.1.2 Full-page Parsing Data Construction
Poster and Chart. The poster data features sparse textual content, making it straightforward to grasp the dependencies between text blocks. Therefore, we organize the paragraph-level text and their corresponding bounding boxes in a left-to-right, top-to-bottom scanning order to create the full-page parsing annotation in JSON format. In contrast, given the highly organized nature of the chart data, we structure its full-page parsing annotations as a JSON dictionary containing hierarchically organized components: chart title, axis labels, legends, and data markers—each precisely associated with their spatial coordinates.
PDF Document. PDF document contains extensive text and complex layouts, necessitating parsing in a logical reading order to comprehend the information accurately. MinerU [52] contains a layout detection model and an OCR model, allowing for extracting text blocks with a certain reading order. However, it often fails to capture all content in documents, missing tables, footnotes, and other elements. In contrast, PyMuPDF [11] can thoroughly extract content from documents, but it does not provide an appropriate reading order. Therefore, we propose a Merge Strategy that combines the strengths of MinerU and PyMuPDF to achieve comprehensive, layout-aware annotations. As shown in Fig. 3(b), 1) we first compare text blocks from ordered and unordered maps to eliminate duplicate text blocks. For the truncated blocks (red block 1 and red block 2 in the ordered map), we replace them with the corresponding complete block (blue block in the unordered map) to improve the semantic completeness within blocks; 2) In the ordered map, we construct an ordered area if two consecutive blocks are placed from the top left to the bottom right. We classify the preserved blocks in the unordered map into two categories: in-area blocks and out-of-area blocks; 3) We insert the in-area blocks into the ordered area and sequentially update the block order in each ordered area using column-major order; 4) For the out-of-area blocks, we insert them into the ordered map. The order of these blocks is then determined based on their positional relationship with their nearest ordered blocks, following column-major order.
@@ 3.2 Instruction-Tuning Data Construction
To seamlessly integrate grounding and referring capabilities in the dialogue and reasoning, it is crucial to construct high-quality instruction-tuning data with accurate grounded text in both query and response. Based on the precise bounding box annotations from Sec. 3.1, we leverage GPT-4o [16] to generate diverse-formatted instruction-tuning data tailored to various text granularities, encompassing tasks such as question answering, reasoning, and summarization. To further enhance instruction-following quality, we first constrain the length and type of the generated responses, resulting in different response categories (e.g., short responses, open-ended long responses). Then, we incorporate a response format prompt into each query. The detailed prompts for GPT-4o are shown in Appendix B
Poster and Chart. As introduced in Sec. 3.1.2, the full-page parsing annotation for poster and chart data includes all text blocks along with their corresponding bounding boxes, which contain sufficient information to construct ground-and-refer instruction-tuning data. Therefore, we provide GPT-4o with the full-page parsing data, and task GPT-4o to generate queries and responses in a grounded manner, where the generated content must include texts that originate from the given parsing annotations. Additionally, we require GPT-4o to wrap the texts originating from parsing annotations with ‘<ocr></ocr>’ and append the corresponding coordinates wrapped in ‘<bbox></bbox>’.
PDF Document. PDF documents often contain extensive text content. If we input all the text of various granularities into GPT-4o, it would result in excessively long prompts. This not only incurs significant API calling costs but also increases the difficulty for GPT-4o in comprehending and completing the tasks, leading to low generation quality. To address these issues, we design a Post-annotating Strategy to generate high-quality data with minimal overhead. Specifically, 1) We first use the document image as input for GPT-4o instead of the text of the document, which significantly reduces the token count; 2) Then, we task GPT-4o to generate queries and responses based on the document images. Additionally, any text within the queries and responses that originates from the document image must be wrapped with “<ocr></ocr>”; 3) We extract the texts wrapped with “<ocr></ocr>” and utilize PyMuPDF to retrieve the corresponding bounding boxes; 4) Finally, we wrap these bounding boxes with ‘<bbox></bbox>’ and insert them back into the generated content after their corresponding texts. For texts that cannot be located by PyMuPDF, we remove wrapped tokens and convert them to plain text. This approach allows the generation of diverse instruction-tuning data at a low cost while ensuring the quality of grounding and referring annotations.